# Rel-A.I.: An Interaction-Centered Approach To Measuring Human-LM Reliance

### v1

@@ 4 Influence of Prior Interactions (RQ1)
The first key shift in human-LM interactions is the frequency in which LMs are being used in everyday human tasks, with recent work showing that over 40% of survey participants were using chat models on a daily basis Wang et al. (2024). Humans are known to develop mental models of systems Norman (1988) and the AI agents Kulesza et al. (2012); Gero et al. (2020) they interact with and update these mental models as interactions progress Bansal et al. (2019). Prior work (see §2) on calibration has made the implicit assumption that human-LM interactions can be evaluated independently from one another. As long as each individual expression is calibrated with the right human behavior, then the model is considered correctly calibrated.
@@
Our first experiment asks how differences in prior interactions, specifically differences in an agent’s verbalized confidence level, might influence human reliance decisions. Each user interacts with two agents, $S_{1A_{conf}}$ and $S_{1B_{unconf}}$, to answer 30 identical geography trivia questions. $S_{1A_{conf}}$ uses only statements of high confidence (strengtheners) and moderate confidence (weakenened strengtheners). $S_{1B_{unconf}}$ uses only statements of moderate confidence (weakenened strengtheners) and low confidence (weakeners). We measure if the reliance rate on these otherwise identical weakened strengtheners (e.g., “I’m pretty sure it’s…”) for these two systems. We hypothesize that weakened strengtheners, which are ambiguously interpreted by users Budescu et al. (1988); Chesley (1986) and most commonly generated by LMs Zhou et al. (2024), are more susceptible to changes in interpretation.
Epistemic Marker $S_{1A-conf}$ $S_{1B-unconf}$
@@ Results
Statements of moderate certainty are relied on significantly less in a confident model (as verbalized) than in an unconfident model (49% compared to 54%, respectively; see Table 2). Specifically, we see expressions like “I’m pretty sure it’s…”, “I think it’s…”, and “It’s likely it’s…”, relied upon 20%, 18%, and 15% more frequently in $S_{1B_{unconf}}$ than in $S_{1A_{conf}}$. In sum, a moderate expression appearing from a confident system will be relied on less than the same expression coming from a less confident system.
We also observe that the effect of increased reliance is not uniform across all templates. Expressions that are the least relied-upon in $S_{1A_{conf}}$ showed the greatest increase in reliance, up to 20%. Reliance on $S_{1A_{conf}}$ and the reliance difference between $S_{1A_{conf}}$ and $S_{1B_{unconf}}$ is negatively correlated with Pearson’s $\rho=-0.38$, see Figure 3.
At the perception level, we see that $S_{1B_{unconf}}$ is generally seen as less competent (2.45 out of 5) than $S_{1A_{conf}}$ (3.35 out of 5) and that humans are also significantly less willing to work with $S_{1B_{unconf}}$ (39.8% vs 74%).
Figure 3: Reliance of expressions from $S_{1A_{conf}}$ versus from $S_{1B_{unconf}}$ (not confident). The less frequently the expressions were relied on in the $S_{1A_{conf}}$, the greater difference between $S_{1B_{unconf}}$ $S_{1A_{conf}}$.
@@ Discussion
These results illustrate how expressions of (un)certainty can shift in meaning depending on the distribution of confidence they appear in. This is consistent with prior work in psycholinguistics by Schuster and Degen (2019) which shows that humans adapt quickly the probabilities they assign to other (human) speaker’s use of epistemic markers. This is particularly worrisome as canonical metrics of machine learning and linguistic calibration do not have ways to account for this critical contextual cue. One could imagine $S_{1B_{unconf}}$ and $S_{1A_{conf}}$ both as “perfectly calibrated” models in the machine learning and linguistic sense. How humans will actually rely on these statements, however, might not reflect this claim.
5 Varying Degree of Warmth (RQ2)
Another key reconfiguration of human-LM interactions has been the anthropomorphism introduced in language technologies Abercrombie et al. (2023b); Araujo (2018); Colombatto and Fleming (2024). Anthropomorphic system attributions like names Wagner et al. (2019) as well as human-like language such as greetings and expressions of warmth Bai et al. (2022) are coveted features as they enable more engaging and enriching human-model interaction. However, the very same anthropomorphic tactics, albeit unintentional, hold the potential for encouraging users to unquestioningly rely on system outputs that are prone to errors and hallucinations Cheng et al. (2024); Abercrombie et al. (2023a). In this section, we investigate how the rise of features of anthropomorphism impact human reliance in decision-making.
Experimental Set-Up
In this experimental setting, each user interacts with three agents to answer 30 geography questions. All the agents answer the same questions and use the exact same distribution of epistemic markers (for distribution of epistemic markers, see Table 1). However, we perturb the presentation of each agent to include two variations of warmth attributes. Our first agent, $S_{2A}$, is given an anonymous name, Agent37821.
Recognizing that publicly deployed language models are given names like Claude and Gemini, we are curious in how the presentation of language models via their name might impact human reliance behaviors. Thus, we create a second agent, $S_{2B}$, which has a humanlike name, Gordon.
Since publicly deployed models have been tuned to incorporate helpfulness Bai et al. (2022), we are interested in the impact on warmth statements on reliance. Leading to our third agent, $S_{2C}$, named Marvin, which will randomly start half of its generations with the phrase, “I’m happy to help!”.
$S_{2A}$ and $S_{2B}$ allows for comparison between anonymous and humanlike names meanwhile $S_{2B}$ and $S_{2C}$ allows for comparison between systems that use warmth cues and systems that don’t. Users interact with all three systems at once in RQ2, allowing us to compare reliance across two dimensions of anthropomorphism.
$S_{2A}$ $S_{2B}$ $S_{2C}$
Warmth 5-point Likert scale 2.533 2.682 3.458*
Competence 2.945 3.046 3.067
Humanlike 2.482 2.661 2.903
Willing to Work Again 46.6% 47.6% 52.3%
Table 3: Perceptions of warmth, competence, humanlikeness, and willingness to work again. $S_{2C}$’s perceived warmth is significantly higher than that of the other two systems; however, we see comparable ratings for competence and willingness to work again.
$S_{2A}$ $S_{2B}$ $S_{2C}$ ${S_{2C}}_{happy}$
Strength. 86.4 88.0* 89.20* 90.98*
Weak. Str. 47.4 49.3* 52.7* 57.37*
Weakener 6.4 9.6* 8.0 9.42
Table 4: Reliance rate (%) across the three systems based on degrees of warmth. ${S_{2C}}_{happy}$ indicates the phrases in $S_{2C}$ that began with a greeting (i.e., “I’m happy to help!”) which were the most relied on expressions. *Significantly higher than the value left of it under bootstrap resampling and two-sampled $t$-test.
Results
The shift from using an anonymous name to a humanlike name does not present significant changes to reliance or perceptions of warmth and competence, consistent with prior work Liu et al. (2023b). However, the introduction of the occasional greetings or politeness (i.e., “I’m happy to help”) is sufficient to significantly increase the perceived warmth of the whole system (Table 3). With this increase in warmth, we also see increases in competence, humanlikeness, and willingness to work with the system again. This suggests that humans are influenced by cues of warmth expressed in the language, and perhaps less so by the humanlike (anthropomorphic attributes) ascribed to the system.
Additionally, as human perceived warmth increases, so does reliance on expressions of (un)certainty. Specifically, we see that participants rely on $S_{2C}$ statements of both strengtheners and weakened strengtheners more frequently than in the two other systems (91% vs 89% and 87%, 57% vs 44% and 47%). Statements with the “I’m happy to help” attribute (${S_{2C}}_{happy}$), were relied on the most (Table 4). Interestingly, among weakeners, $S_{2B}$ was relied upon the most across all three systems, illustrating that the effect of warmth isn’t necessarily uniform across all degrees of certainty. Within perturbations of names, we see that $S_{2B}$ was relied on significantly more often than $S_{2A}$. Lastly, there exists a strong correlation between perceived warmth and reliance. Aggregating across all three systems, as perceived warmth increases, reliance also increases from 35% to 64% (Pearson’s $\rho=0.97$, Figure 4).
Figure 4: Correlation of reliance and self-reported perception of warmth, aggregated across all three models Pearson’s $\rho=0.97$.
Discussion
Work from Cheng et al. (2022) illustrates that the anthropomorphism of chat models in consumer settings leads to increased perceptions of trust, Kim et al. (2024) shows in a medical setting that hedges with personal pronouns are less likely to be relied upon than those without, meanwhile Inie et al. (2024) illustrates the interaction effects between anthropomorphism and product type. Our work adds to prior work and shows that warmth, a product of anthropomorphic cues, impacts user interpretation at a system-level, suggesting a more careful use of these cues may be warranted.
Lastly, in human perceptions of humans, warmth and competence have a hydraulic relationship where when perceptions of one goes up, the other goes down Cuddy et al. (2011). However, our results show that as perceived warmth increases, perceived competence does not decrease, in fact — we see a slight increase (Table 3). These findings potentially highlight that the hydraulic relationship of warmth and competence may apply differently to perceptions of AI agents.
@@ 6 Domain Specific Reliance (RQ3)
math abstract algebra philo- sophy world religion law
@@ Discussion
Beyond verbalized confidences of LMs, exogenous factors like subject domain influence human reliance on LMs. These findings are consistent with results from Cohn et al. (2024), where they found that human ratings of generations in medical domains differed from those of generations in career domains.
One potential hypothesis for this difference may be due to how humans view AI agents as machines proficient in calculations. Ironically, in the age of LMs, chat models aren’t actually performing computations under the hood but rather are simply completing the given prompt. As language technology advances, the mental models of prior technologies might still be influencing human decision making and not all domains should be treated equally in interaction calibration.
@@ 7 Discussion and Conclusion
In this work, we introduce Rel-A.I., an interaction-centered approach for evaluating human reliance on LM-generated responses. With this methodology, we tackle three emergent properties of human-LM interactions and ask how these new characteristics might influence human decision-making. We find that prior interactions, the warmth of the model, and the subject domain of the interaction can all impact how and when humans choose to rely on the exact same expressions of (un)certainty. These findings illustrate the limitations of prior in vitro evaluations which exclusively focused on the calibration between internal model probabilities and the verbalized confidence of LMs. We now discuss the implications and aspirations of our contributions.
@@
We must reorient ourselves away from model calibration and instead look towards calibrating human-LM interactions. Although perfectly (linguistically) calibrated language models could exist, these models might not actually yield safe and optimal human-LM interactions. Humans will interpret the meaning of verbalized confidences and these interpretations will consider the interaction context and the users’ history, adding exogenous features to how a user might actually behave towards LM-generated epistemic markers.
The Cost of Warmth Cues
With the rise of anthropomorphism and friendliness of models, there is a trend to build more engaging and personable LMs. However, our findings illustrate that these new dimensions could potentially introduce unexpected safety risks. Our findings show that warmth and reliance are tightly coupled and the introduction of a seemingly benign friendly greeting could actually encourage a user to rely on the system more. As designers and practitioners think of ways to build personable and engaging interactions between humans and LMs, they must also consider the safety implications these decisions have on human overreliance.
@@ A.1 Recruitment Process Details
A.2 Template Elicitation Details
Following prior work Zhou et al. (2024), we elicited expressions of uncertainty from nine publicly deployed language models (text-davinci-003, GPT-3.5-Turbo, GPT-4, LLaMA-2 7B, LLaMA-2 13B, LLaMA-2 70B, Claude-1, Claude-2, Claude- Instant-1) and performed qualitative coding Auerbach and Silverstein (2003); Charmaz (2006) to code the expressions The most frequently LM-generated expressions were categorized into weakeners (e.g., ”I’m not sure…”, ”Maybe it’s…”), strengtheners (e.g., ”I’m sure it’s…”, ”Undoubtedly it’s…”),
A.3 Details on Experiments on Anthropomorphism
@@ A.4 Pilot Experiment for RQ3
A.5 Screenshots of User Tasks
@@
Figure 7: Example Task

### v2

@@ 5 Influence of Prior Interactions (RQ2)
The next shift in human-LM interactions is the frequency in which LLMs are being used in everyday human tasks, with recent work showing that over 40% of survey participants were using chat models on a daily basis Wang et al. (2024). Humans are known to develop mental models of systems Norman (1988) and the AI agents Kulesza et al. (2012); Gero et al. (2020) they interact with and update these mental models as interactions progress Bansal et al. (2019). Prior work (see §2) on calibration has made the implicit assumption that human-LM interactions can be evaluated independently from one another. As long as each individual expression is calibrated with the right human behavior, then the model is considered correctly calibrated.
@@
Our experiment asks how differences in prior interactions, specifically differences in an agent’s verbalized confidence level, might influence human reliance decisions. Each user interacts with two agents, $B_{conf}$ and $B_{unconf}$, to answer 30 identical geography trivia questions. $B_{conf}$ uses only statements of high confidence (strengtheners) and moderate confidence (weakenened strengtheners). $B_{unconf}$ uses only statements of moderate confidence (weakenened strengtheners) and low confidence (weakeners). We measure if the reliance rate on these otherwise identical weakened strengtheners (e.g., “I’m pretty sure it’s…”) for these two systems. We hypothesize that weakened strengtheners, which are ambiguously interpreted by users Budescu et al. (1988); Chesley (1986) and most commonly generated by LLMs Zhou et al. (2024), are more susceptible to changes in interpretation.
Epistemic Marker $B_{conf}$ $B_{unconf}$
@@ Results
Statements of moderate certainty are relied on significantly less in a confident model (as verbalized) than in an unconfident model (49% compared to 54%, respectively; see Table 3). Specifically, we see expressions like “I’m pretty sure it’s…”, “I think it’s…”, and “It’s likely it’s…”, relied upon 20%, 18%, and 15% more frequently in $B_{unconf}$ than in $B_{conf}$. In sum, a moderate expression appearing from a confident system will be relied on less than the same expression coming from a less confident system.
We also observe that the effect of increased reliance is not uniform across all templates. Expressions that are the least relied-upon in $B_{conf}$ showed the greatest increase in reliance, up to 20%. Reliance on $B_{conf}$ and the reliance difference between $B_{conf}$ and $B_{unconf}$ is negatively correlated with Pearson’s $\rho=-0.38$, see Figure 4.
At the perception level, we see that $B_{unconf}$ is generally seen as less competent (2.45 out of 5) than $B_{conf}$ (3.35 out of 5) and that humans are also significantly less willing to work with $B_{unconf}$ (39.8% vs 74%).
Figure 4: Reliance of expressions from $B_{conf}$ versus from $B_{unconf}$ (not confident). The less frequently the expressions were relied on in the $B_{conf}$, the greater the difference between $B_{unconf}$ and $B_{conf}$.
@@ Discussion
These results illustrate how expressions of (un)certainty can shift in meaning depending on the distribution of confidence they appear in. This is consistent with prior work in psycholinguistics by Schuster and Degen (2019) which shows that humans adapt quickly the probabilities they assign to other (human) speaker’s use of epistemic markers. This is particularly worrisome as canonical metrics of machine learning and linguistic calibration do not have ways to account for this critical contextual cue. One could imagine $B_{unconf}$ and $B_{conf}$ both as “perfectly calibrated” models in the machine learning and linguistic sense. How humans will actually rely on these statements, however, might not reflect this claim.
@@ 6 Domain Specific Reliance (RQ3)
math algebra philos. religion law
@@ Discussion
Beyond verbalized confidences of LLMs, exogenous factors like subject domain influence human reliance on LLMs. These findings are consistent with results from Cohn et al. (2024), where they found that human ratings of generations in medical domains differed from those of generations in career domains.
One potential hypothesis for this difference may be due to how humans view AI agents as machines proficient in calculations. Ironically, in the age of LLMs, chat models aren’t actually performing computations under the hood but rather are simply completing the given prompt. As language technology advances, the mental models of prior technologies might still be influencing human decision making and not all domains should be treated equally in interaction calibration.
@@ 7 Discussion and Conclusion
In this work, we introduce Rel-A.I., an interaction-centered approach for evaluating human reliance on LM-generated responses. With this methodology, we tackle three emergent properties of human-LM interactions and ask how these new characteristics might influence human decision-making. We find that prior interactions, the warmth of the model, and the subject domain of the interaction can all impact how and when humans choose to rely on the exact same expressions of (un)certainty. These findings illustrate the limitations of prior in vitro evaluations which exclusively focused on the calibration between internal model probabilities and the verbalized confidence of LLMs. We now discuss the implications and aspirations of our contributions.
@@
We must reorient ourselves away from evaluations on language quality along and instead look towards evaluating human-LM interactions. Perfectly (linguistically) calibrated language models might actually yield unexpected human behavior responses if interactional contexts are not carefully considered.
The Cost of Greetings
With the rise of anthropomorphism and friendliness of models, there is a trend to build more engaging and personable LLMs. However, our findings illustrate that these new dimensions could potentially introduce unexpected risks. Perceived competence can easily be altered and has significant consequence on human reliance rates. As designers and practitioners think of ways to build personable and engaging interactions between humans and LLMs, they must also consider the implications these decisions have on human overreliance.
@@ A.1 Recruitment Process Details
A.2 Details on Experiments on Anthropomorphism
@@ A.3 Pilot Experiment for RQ3
A.4 Screenshots of User Tasks
@@
Figure 7: Example task where users decide whether or not to rely on an agent’s response or to look up the answer themselves.
A.5 Perceived Competence as a Significant Variable for Explaining Variance in Human Reliance
Since each greeting can affect multiple dimensions of perception, we perform ordinary least squares analysis to disentangle the effects of perceived competence, warmth, and humanlikeness explain the variance on reliance. Perceived competence, warmth, and humanlikeness explain 33%, 7%, and 6% of the variance in reliance respectively, each with significant positive coefficients. With the addition of the random variable, UserId, the four variables explain 34% of the variance in reliance with perceived competence and UserId having significant positive coefficient ($p<0.05$). See tables 8, 9, 10, and 11 for details.
Dep. Variable: rely R-squared: 0.330
Model: OLS Adj. R-squared: 0.329
Method: Least Squares F-statistic: 367.9
Date: Thu, 19 Sep 2024 Prob (F-statistic): 5.40e-67
Time: 13:48:34 Log-Likelihood: 152.08
No. Observations: 750 AIC: -300.2
Df Residuals: 748 BIC: -290.9
Df Model: 1
Covariance Type: nonrobust
coef std err t P>|>|> |t|||| [0.025 0.975]
const 0.0892 0.024 3.792 0.000 0.043 0.135
competent 0.1469 0.008 19.181 0.000 0.132 0.162
Omnibus: 1.453 Durbin-Watson: 1.051
Prob(Omnibus): 0.484 Jarque-Bera (JB): 1.324
Skew: -0.043 Prob(JB): 0.516
Kurtosis: 3.187 Cond. No. 11.0
Table 8: OLS Regression Results with perceived competence as dependent variable and reliance as independent variable.
Dep. Variable: rely R-squared: 0.072
Model: OLS Adj. R-squared: 0.071
Method: Least Squares F-statistic: 58.23
Date: Thu, 19 Sep 2024 Prob (F-statistic): 7.12e-14
Time: 13:48:34 Log-Likelihood: 30.181
No. Observations: 750 AIC: -56.36
Df Residuals: 748 BIC: -47.12
Df Model: 1
Covariance Type: nonrobust
coef std err t P>|>|> |t|||| [0.025 0.975]
const 0.3506 0.024 14.871 0.000 0.304 0.397
warmth 0.0593 0.008 7.631 0.000 0.044 0.075
Omnibus: 20.819 Durbin-Watson: 0.950
Prob(Omnibus): 0.000 Jarque-Bera (JB): 10.714
Skew: -0.044 Prob(JB): 0.00472
Kurtosis: 2.421 Cond. No. 9.22
Table 9: OLS Regression Results with perceived warmth as dependent variable and reliance as independent variable.
Dep. Variable: rely R-squared: 0.061
Model: OLS Adj. R-squared: 0.060
Method: Least Squares F-statistic: 48.54
Date: Thu, 19 Sep 2024 Prob (F-statistic): 7.09e-12
Time: 13:48:34 Log-Likelihood: 25.650
No. Observations: 750 AIC: -47.30
Df Residuals: 748 BIC: -38.06
Df Model: 1
Covariance Type: nonrobust
coef std err t P>|>|> |t|||| [0.025 0.975]
const 0.3708 0.023 16.225 0.000 0.326 0.416
humanlike 0.0537 0.008 6.967 0.000 0.039 0.069
Omnibus: 42.609 Durbin-Watson: 0.911
Prob(Omnibus): 0.000 Jarque-Bera (JB): 16.817
Skew: -0.055 Prob(JB): 0.000223
Kurtosis: 2.275 Cond. No. 8.71
Table 10: OLS Regression Results with perceived humanlikeness as dependent variable and reliance as independent variable.
Dep. Variable: rely R-squared: 0.335
Model: OLS Adj. R-squared: 0.331
Method: Least Squares F-statistic: 93.62
Date: Thu, 19 Sep 2024 Prob (F-statistic): 1.66e-64
Time: 13:48:34 Log-Likelihood: 154.78
No. Observations: 750 AIC: -299.6
Df Residuals: 745 BIC: -276.5
Df Model: 4
Covariance Type: nonrobust
coef std err t P>|>|> |t|||| [0.025 0.975]
const 0.1048 0.026 4.059 0.000 0.054 0.156
competent 0.1502 0.009 16.417 0.000 0.132 0.168
warmth -0.0081 0.009 -0.957 0.339 -0.025 0.009
humanlike 0.0008 0.008 0.105 0.917 -0.015 0.017
random_id -0.1971 0.094 -2.105 0.036 -0.381 -0.013
Omnibus: 1.759 Durbin-Watson: 1.053
Prob(Omnibus): 0.415 Jarque-Bera (JB): 1.642
Skew: -0.057 Prob(JB): 0.440
Kurtosis: 3.199 Cond. No. 67.8
Table 11: OLS Regression Results with perceived competence, warmth, humanlikeness as dependent variables, random_id as a random user variable, and reliance as independent variable.